OpenStudy (anonymous):

find the derivative of the function y=1n 3sqrt(x-1/x+1 (fan & award best answer)

OpenStudy (abb0t):

What?

OpenStudy (abb0t):

What??????????

OpenStudy (anonymous):

thanks !! [:

OpenStudy (fibonaccichick666):

np, do you know how to solve it from here?

OpenStudy (anonymous):

i think s im working on a another problem and come back to it

OpenStudy (anonymous):

@FibonacciChick666 no i dont know how to finish it

OpenStudy (abb0t):

Why would you use u-sub? You need to know derivatives before you use u-sub for integration

OpenStudy (abb0t):

You can apply your log rules from algebra to split everything up

OpenStudy (anonymous):

I believe that 3 square root should be read as the cube root from what I have seen from other posters.

OpenStudy (abb0t):

ln(ab) = ln(a) + ln(b)

OpenStudy (anonymous):

And if so, \[y=\ln \sqrt[3]{\frac{ x-1 }{ x+1 }} \] can be simplified to by rules of natural logs to \[y=\frac{ 1 }{ 3 }\ln \frac{ x-1 }{ x+1 }\] which can then be further simplified by natural logs to \[y=\frac{ 1 }{ 3 }\ln(x-1)-\frac{ 1 }{ 3 }\ln(x+1)\] Now, if we take the derivative of y with respect to x, we get \[\frac{ dy }{ dx }=\frac{ 1 }{ 3(x-1) }-\frac{ 1 }{ 3(x+1) }\] This can be further simplified by algebra to: \[\frac{ x+1-(x-1) }{ 3(x+1)(x-1) }\] Finally, combining like terms in the numerator, we get: \[\frac{ 2 }{ 3(x+1)(x-1) }\]

OpenStudy (abb0t):

Follow @calmat01 method. He has it down.

OpenStudy (fibonaccichick666):

ok so how to solve this problem: (I disagree with calmat01's answer.) First let me restate the question: \[y=ln(3((x−1)/(x+1))^{.5})\] Now first rule of natural logs. You can take the exponent down and multiply it in front. So we have: \[y=0.5*ln(3((x−1)/(x+1))\] Second rule of logs. A quotient can be rewritten as a subtraction. So we have: \[y=0.5*[ln(3(x−1))-ln(x+1)]\] Next we do u,v substitution. let u=3x-3 du=? v=x+1 dv=? Can you please compute these derivatives and see if you can finish the problem from this step? \[y=0.5*[ln((u)-ln(v)]\]

OpenStudy (abb0t):

You basically did the same thing as calmat except you used substitution which I think makes it more confusing when computing derivatives of natural log functions.

OpenStudy (fibonaccichick666):

You are required to use substitution because it is not of the standard form. You also do not make the 3 a cube root; therefore the answer is wrong.

OpenStudy (fibonaccichick666):

then again, I also interpreted the original question as 3 times the square root not a cube root.

OpenStudy (anonymous):

I was basing my approach to the problem by what I have seen using the equation editor. In the equation editor, problems written in the form a root x become \[\sqrt[a]{x}\] which is how I interpreted his question.